OpenStudy (anonymous):

Partial fraction decomp: 1/(s^2+4)(s^2+1)=???

OpenStudy (anonymous):

can this be decomposed?

hartnn (hartnn):

if you want it in the form of \(\large \dfrac{A}{s^2+4}+\dfrac{B}{s^2+1}\) then, yes. i would suggest you consider s^2 as x and then do an easy partial fraction for, \(\dfrac{1}{(x+4)(x+1)}\)

OpenStudy (anonymous):

Lap^-1[e^-2PI*s/(s^2+4)(s^2+1)]

OpenStudy (anonymous):

hmm okay

hartnn (hartnn):

so, yes, you need it in that form, can you do the partial with x that i mentioned ?

OpenStudy (anonymous):

yeah pretty sure. hold on

OpenStudy (anonymous):

(1/3)*1/(s^2+1)+(1/5)*1/(s^2+4)

OpenStudy (anonymous):

okay, so -1/3... no 1/5. Got it thank you!

OpenStudy (anonymous):

I didn't think to replace s^2 with x.. That took care of it.

hartnn (hartnn):

your welcome ^_^